OpenStudy (anonymous):

Find the vertices of the hyperbola with equation: (x+2)^2/4 - (y+1)^2/16. very last question!

OpenStudy (amistre64):

yay!! the vertices are along the + variable; and that is.....x axis

OpenStudy (amistre64):

the major line is centered at (-2,-1) so the verticeses are (+-x, -1)

OpenStudy (amistre64):

make y=0 to find x :)

OpenStudy (amistre64):

x+2 = +-2 x = 0 and -4

OpenStudy (amistre64):

(0,-1) and (-4,-1)

OpenStudy (amistre64):

im pretty sure thats it

OpenStudy (anonymous):

ok that is what I got too so i think that it is correct :)

OpenStudy (amistre64):

-2+2 = 0 -2-2 = -4; its good :)